OpenStudy (anonymous):

How many terms of the series -9,-6,-3..... must be taken so that the sum may be 66

OpenStudy (anonymous):

I reached a quadratic that can't be simplified

ganeshie8 (ganeshie8):

they all are negative numbers right ?

ganeshie8 (ganeshie8):

how can u get a positive sum ?

OpenStudy (anonymous):

the series continues , it is in A.P with a common difference "-3"

ganeshie8 (ganeshie8):

Oh yes, my brain is 100% dead today :o

OpenStudy (anonymous):

loolol happens

ganeshie8 (ganeshie8):

so are we trying to solve : \[\large \dfrac{n}{2}[2(-9) + (n-1)(3)] >= 66\] ?

ganeshie8 (ganeshie8):

d = -6-(-9) = -6+9 = 3

OpenStudy (anonymous):

oh then i can do it now , i took d as -3

ganeshie8 (ganeshie8):

equality will do i think...

OpenStudy (anonymous):

So i am getting n^2 -7n +44 = 0

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

should get n^2 -7n \(\color{Red}{-}\) 44 = 0

OpenStudy (anonymous):

Yeah sry , yes

ganeshie8 (ganeshie8):

-11, 4 will do